### v1

# Whistle: Data-Efficient Multilingual and Crosslingual Speech Recognition via Weakly Phonetic Supervision

Abstract
There exist three approaches for multilingual and crosslingual automatic speech recognition (MCL-ASR) - supervised pre-training with phonetic or graphemic transcription, and self-supervised pre-training. We find that pre-training with phonetic supervision has been underappreciated so far for MCL-ASR, while conceptually it is more advantageous for information sharing between different languages. This paper explores the approach of pre-training with weakly phonetic supervision towards data-efficient MCL-ASR, which is called Whistle. We relax the requirement of gold-standard human-validated phonetic transcripts, and obtain International Phonetic Alphabet (IPA) based transcription by leveraging the LanguageNet grapheme-to-phoneme (G2P) models. We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, with 10 seen languages and 2 unseen languages. A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup for MCL-ASR. Experiments demonstrate the advantages of phoneme-based models (Whistle) for MCL-ASR, in terms of speech recognition for seen languages, crosslingual performance for unseen languages with different amounts of few-shot data, overcoming catastrophic forgetting, and training efficiency. It is found that when training data is more limited, phoneme supervision can achieve better results compared to subword supervision and self-supervision, thereby providing higher data-efficiency. To support reproducibility and promote future research along this direction, we will release the code, models and data for the whole pipeline of Whistle at https://github.com/thu-spmi/CAT upon publication.
Index Terms:
speech recognition, multilingual, crosslingual, data-efficient, IPA.
I Introduction
In recent years, deep neural network (DNN) based automatic speech recognition (ASR) systems have achieved significant progress, which are, however, data-hungry. A substantial amount of transcribed speech data are required for model training. There are more than 7,000 languages spoken around the world [1], but due to the lack of training data, only a small fraction of them benefit from current ASR technology. An important challenge for the speech community is that we can develop ASR systems to new unsupported languages rapidly and at reasonable costs. Multilingual and crosslingual ASR (MCL-ASR) have been studied as an effective way to address this problem.
In multilingual speech recognition, training data for a number of languages, often referred to as seen languages, are merged to train a multilingual model, which can be used to recognize speech from all seen languages. The multilingual model can also serve as a pre-trained model, which can be further fine-tuned for crosslingual speech recognition. Crosslingual speech recognition refers to recognizing utterances in a new language, which is unseen in training the multilingual model. From machine learning perspective, such multilingual and crosslingual training can be regarded as performing multi-task learning and transfer learning, which promotes sharing of statistical strength. The advantage is that the ASR performance for low-resource languages, both seen and unseen, can be improved, and the cost of system building and maintenance for multiple languages can be reduced as well.
The general concept of multilingual and crosslingual speech recognition has been applied for a long time, dating back to the time when GMM-HMM based classic models and then DNN-HMM based hybrid models are prevalent in ASR research, to name a few, e.g., in [2] and [3] respectively. Recently, end-to-end models have emerged [4, 5, 6], which can be directly trained from phonetic or graphemic transcription, eliminating the first pass of producing HMM state alignment. For end-to-end models, the approach of pre-training followed by fine-tuning has attracted increasing interests and achieved good performance. There are mainly two classes of pre-training methods, based on either self-supervised learning or supervised learning. Self-supervised pre-training is conducted over unlabeled speech data from multiple languages for speech representation learning in general [7, 8, 9]. Supervised pre-training, by applying end-to-end models on multilingual labeled speech data, can be further divided into two sub-categories of research, which are contrasted by using different types of modeling units. The first is grapheme-based or subword-based [10, 11, 12, 13], which, collectively referred to as based on graphemic transcription (orthography), creates a shared token set across multiple languages, e.g., using 10K sentence pieces [11]. The second trains end-to-end models on phonetic transcriptions [14, 15, 16, 17, 18], which usually utilizes International Phonetic Alphabet (IPA) symbols to create a (nearly-)universal phone inventory, e.g., using 187 phones [14].
Intuitively, the key to successful multilingual and crosslingual recognition is to optimize information sharing during multilingual training and maximize the knowledge transferring from a well trained multilingual model to the model trained for recognizing utterances in a new language [15]. Taking this perspective, we could examine the pros and cons of the three approaches - supervised pre-training with graphemic transcription or phonetic transcription, and self-supervised pre-training, which is detailed in Section II.
While requiring pronunciation lexicons, pre-training with phonetic supervision is more advantageous for information sharing between different languages. For phonetic supervision, IPA symbols include enough symbols to represent the fundamental sounds of all languages, and sounds in different languages share these phonetic representations [19]. In contrast, graphemes and subwords are in fact from writing systems of languages (orthography), not for describing and distinguishing all the sounds in human language throughout the world, which is exactly phonetic transcription does. Creating a graphemic token set from multiple languages for supervision is non-trivial and delicately affects ASR performance; until recently, tokenization strategy is still under investigation and needs a balance between granularity and ASR performance [12]; adding new languages for crosslingual recognition further complicates the design of tokenization. Besides the above theoretical analysis of supervised pre-training with graphemic transcription and phonetic transcription, an interesting research question is about empirical comparison. It has been empirically found that compared to learning with graphemic supervision, learning with phonetic supervision performs equally strong and tends to be more data-efficient in monolingual ASR [20, 21, 22, 23]. But to the best of our knowledge, there have been no solid experiments to study which approach is better or if they yields similar results for MCL-ASR, when evaluated in a common experimental setup (Research Question 1, referred to as RQ-1).
To address the problem of requiring phonetic transcription for phonetic supervision, we note that phonetic resources and tools have been steadily developed over these years and are easily accessible, including grapheme-to-phoneme (G2P) models and tools [24, 25, 26], phoneme inventories [27]. We can relax the requirement of gold-standard human-validated transcripts, and in this paper, we obtain the IPA phonetic transcripts by leveraging the LanguageNet G2P models [25]. The LanguageNet G2P models are available for 142 languages, with the phoneme error rates (PERs) ranging from 7% to 45%. So the main aim of this paper is to investigate weakly supervised pre-training with somewhat noisy phonetic transcription. This is in spirit similar to the work in Whisper [13]. But instead of using weakly graphemic supervision in Whisper, our work employs weakly phonetic supervision. We call the approach investigated in this paper: Whistle (Weakly phonetic supervision strategy for multilingual and crosslingual speech recognition).
A secondary interesting research question is to compare supervised pre-training and self-supervised/un-supervised pre-training. Basically, we agree with the comments in [13]. Current pre-trained models for speech such as based on wav2vec 2.0 [28] aim to learn speech representation in general over unlabeled data; They mostly are encoder-only and thus lack an equivalently performant decoder, which requires at least adding a classifier layer and supervised finetuning over labeled data even for seen languages. These comments, presumably, are suited to comparing self-supervision to both graphemic supervision [13] and phonetic supervision (our work). These being said, to the best of our knowledge, there have been no strict experiments to study which approach is better or if they yields similar results for MCL-ASR, when evaluated in equal settings (Research Question 2, referred to as RQ-2).
In summary, this paper explores supervised pre-training with weakly phonetic supervision, towards data-efficient multilingual and crosslingual speech recognition. Our main contributions are as follows.
  • We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, to evaluate multilingual and crosslingual speech recognition, with 10 seen languages and 2 unseen languages, measuring both phoneme error rate (PER) and word error rate (WER). A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup - supervised pre-training with graphemic transcription or weakly phonetic transcription, and self-supervised pre-training for MCL-ASR. These experiments present our effort to answer RQ-1 and RQ-2.
  • We develop Whistle, an approach to data-efficient multilingual and crosslingual speech recognition via weakly phonetic supervision, including the whole pipeline of data processing, model training and testing. Experiments demonstrate the advantages of Whistle for MCL-ASR, in terms of speech recognition for seen languages, crosslingual performance for unseen languages with different amounts of few-shot data, overcoming catastrophic forgetting, and training efficiency.
  • Many prior works on multilingual and crosslingual speech recognition were conducted on internal or proprietary datasets such as GlobalPhone [29] and IARPA Babel [^1] https://www.iarpa.gov/index.php/research-programs/babel, which are not openly-available. We find that supervised pre-training with phonetic supervision has been underappreciated so far for MCL-ASR. To promote future research along this direction, we release the code, models and data for the whole pipeline of Whistle at the following URL: https://github.com/thu-spmi/CAT.
II Related work
II-A MCL-ASR with phonetic supervision
Research in multilingual and cross-lingual ASR has long been motivated by phonetics and has used phonetic supervision, e.g., in [30, 3, 2, 14, 31, 15, 17], to name a few. The major phonetic alphbet in use is the International Phonetic Alphabet (IPA), which includes modified Roman letters and diacritics, by means of which the sounds of all human languages can be represented [19]. So a common practice is to combine the phonetic inventory of all languages to be recognized into a global phoneme set, often based on IPA. Employing phonetic units is, presumably, the most intuitive way to promote information sharing and learn language-universal representations for MCL-ASR. Modeling based on phonetic supervision further allows to pursue finer level of information sharing by decomposing phones into a list of phonological articulatory attributes [32, 15, 17, 33].
To address the problem of requiring phonetic transcription for phonetic supervision, there have been steady efforts to develop phonetic resources and tools. Epitran provides a 61-language rule-based open-source G2P tool [24]; the LanguageNet includes FST (Finite State Transducer) based G2P models in nearly 150 languages [25], and PHOIBLE compiles a database of phone inventories for more than 2000 languages and dialects [27]. Based on these phonetic resources and tools, there has been continuous studies. Base on Epitran G2P, [14] first predicts over a shared phone inventory, and then introduces an allophone layer to map into language-specific phonemes. 11 training languages and 2 unseen languages were used. Based on LanuageNet G2P, monolingual, multilingual and (zero-shot) crosslingual CTC models are trained over 13 languages in [31], with the output layer consisting of IPA symbols. Every modifier symbol is treated as a separate token, and so phonetic token error rates (PTERs) are measured. Compared to monolingual models, it reports major PTER improvements across all 13 languages in the multilingual setup, and stark degradation in the crosslingual systems. The recent studies [14, 31] mainly investigate universal phone recognition. There remains an interesting question, as also raised in [31], whether improvements in error rates would also be observed in downstream metrics such as WER. Another related question is which approach of phonetic and graphemic supervision is better for MCL-ASR (RQ-1), since no comparison is conducted in these recent multilingual studies.
II-B MCL-ASR with graphemic supervision
Graphemic transcription (orthography), as a part of the writing system in a language, does not represent the sounds of a language in a consistent way [19]. In many languages, there is a discrepancy between graphemic transcription and phonetic transcription. With the learning power of deep neural networks, people has begun to build ASR systems with the output layer consisting of graphemic units such as characters [21], subwords [34, 23], or words [35], initially for monolingual ASR and recently applied to MCL-ASR. Using graphemic supervision eliminates the requirement of pronunciation lexicons for different languages and simplifies the pipeline of MCL-ASR. On the other hand, pooling and creating a large set of graphemic tokens from multiple languages brings the label sparsity issue and the resulting MCL-ASR systems tend to be data-hungry, and tokenization scheme is an active research question [36, 12].
Thanks to larger and larger amounts of transcribed speech data and increasingly large neural networks, subword-based supervised pre-training has obtained better and better performance and become a widely adopted strategy in industry to build MCL-ASR systems for increasingly many languages. For example, the Whisper [13] models use the a Byte-Pair Encoding (BPE) text tokenizer and are trained over 680,000 hours cleaned web data by weakly graphemic supervision, capable of recognizing speech from 97 languages. While achieving impressive performance, recent advances in large MCL-ASR models are presumably an effect of scaling power, and it is hard to argue that the good results are not due to having additional data, nor due to the large neural architecture. It remains unclear which approach (phonetic supervision or grapheme supervision) is better when evaluated in an equal experimental setting, or if they produce similar results for MCL-ASR. This paper presents our preliminary effort to answer this question (RQ-1).
II-C MCL-ASR with self-supervision
Self-supervised learning methods mainly refer to some recent learning methods based on contrastive learning such as wav2vec 2.0 [28] or masking prediction such as BERT [37], which can still be regarded as unsupervised learning methods from a classical perspective (no data annotation is required). Therefore, the literature often does not strictly distinguish between unsupervised and self-supervised learning methods in terms of terminology, and we can collectively referred to them as unsupervised learning methods. Self-supervised learning methods such as wav2vec 2.0 [28] have been proposed to learn speech representation in general from multilingual unlabeled speech data. Based on wav2vec 2.0, XLS-R models [8] are trained on unlabeled data from 128 languages. In the recent Massively Multilingual Speech (MMS) project [9], wav2vec 2.0 based models are pre-trained over 1,406 languages, and CTC based multilingual ASR models for 1,107 languages are then fine-tuned using labeled data for each language. Specifically, a linear layer is added on top of pre-trained MMS models which maps to an output vocabulary which is the set of letters in the labeled training data, and is then fine-tuned with the CTC loss.
As commented in [13], while current unsupervised pre-training has improved the quality of audio encoders, the lack of an equivalently high-quality pre-trained decoder is a crucial weakness which limits their usefulness. In the following, we provide a closely related comment. We find that current unsupervised pre-training methods in learning audio encoders such as wav2vec 2.0 does not satisfy the so-called principled unsupervised learning, since “the unsupervised objective may be unrelated to the supervised task of interest” [38]. In contrast, the GPT based unsupervised pre-training method for natural language processing (NLP) tasks is principled, since the supervised objective is the same as (closely related to) the unsupervised objective but only evaluated on a subset of the sequence in NLP [39]. For ASR tasks, these comments favor supervised pre-training (either grapheme-supervision or phonetic supervision) over the current unsupervised pre-training. These being said, remarkably, it has been known in various machine learning tasks that supervised and unsupervised training methods are not mutually exclusive and could be jointly used to define semi-supervised learning, e.g., in image classification [40], speech recognition [41, 42], natural language labeling [43], dialog systems [44]. A complete investigation into semi-supervised learning for ASR is outside the scope of this paper. This paper presents a straightforward empirical comparison between self-supervision and phonetic supervision for MCL-ASR in a common experimental setup (RQ-2).
Figure 1: Illustration of the pre-training and fine-tuning procedures with phonetic supervision, subword supervision, and self-supervision.
III Approach
In this section, we describe the three main classes of pre-training and fine-tuning methods for MCL-ASR, i.e., phoneme-based multilingual supervised pre-training (Section III-A), subword-based multilingual supervised pre-training (Section III-B) and multilingual self-supervised pre-training (Section III-C). Figure 1 shows the differences between the three methods. We can see from Figure 1 that similar neural network architectures can be used for the acoustic encoders in all the three methods, which is good for fair comparison.
The input to the acoustic encoder is usually spectral features, obtained from short-time Fourier transform frame by frame, denoted by $x_{1},\cdots,x_{T}\triangleq x_{1:T}$. In DNN-based ASR, the acoustic encoder could be viewed as a non-linear feature extractor, which hopefully can be trained to extract high-level features (or say, representations), more discriminative than the raw spectral features. The output representations from the acoustic encoder are denoted by $h_{1},\cdots,h_{T}\triangleq h_{1:T}$. A popular neural network architecture for the encoder is Conformer [45], which consists of convolution blocks followed by Conformer blocks.
Given acoustic observations $x_{1:T}$, the task of ASR is to find the most likely labels $y_{1},\cdots\,y_{L}\triangleq y_{1:L}$. Different units can be used for labeling $y_{1:L}$, depending on what transcription is used for labeling, phonetic or graphemic, as shown in Table II. Phonemes and subwords are two widely-used labels for MCL-ASR.
In order to promote information sharing between different languages for MCL-ASR, training data from a number of languages, often referred to as seen languages, can be merged to pre-train a multilingual encoder in a supervised fashion, with labels of $y_{1:L}$ given in the form of either phonemes or subwords. Alternatively, the acoustic encoder could be pre-trained over unlabeled data by some self-supervised method, such as wav2vec 2.0 [28], and then be fine-tuned over labeled data in the form of either phonemes or subwords.
III-A Phoneme-based multilingual supervised pre-training
[…]
$\begin{split}z_{t}&=W^{T}h_{t}\in\mathbb{R}^{|V|+1}\\ P(\pi_{t}=k|x_{1:T})&=\frac{\exp(z_{t}^{k})}{\sum_{j=1}^{|V|+1}\exp(z_{t}^{j})},k=1,\cdots,|V|+1\end{split}$ (1)
where $W\in\mathbb{R}^{(|V|+1)\times D}$ denotes the weight matrix, and we omit the bias vector in describing the linear layer. The un-normalized outputs $z_{t}$ are often called logits, and $z^{k}_{t}$ denotes the logit corresponding to label $k$.
In phoneme-based multilingual supervised pre-training investigated in this paper, which is called Whistle, we take the union of the phoneme inventories from the seen languages to be the alphabet of labels $V_{\text{multi}}$. The $k$-th row vector from the matrix $W$, denoted by $W(k,:)$, could be viewed as the phoneme embedding for phoneme $k$. The logit for phoneme $k$ at frame $t$ is actually an inner product between the phoneme embedding and the representation vector, $z_{t}^{k}=W(k,:)^{T}h_{t}$.
For recognizing speech from a seen language, the pre-trained encoder together with the phoneme embeddings can be directly used without fine-tuning. Specifically, we build a weighted finite state transducer (WFST) [46], obtained by composing the CTC topology, pronunciation lexicon and word-level n-gram language model, and use WFST-based decoding [47, 21]. While requiring pronunciation lexicons (PROLEX), pre-training with phonetic supervision is more advantageous for information sharing between different languages. In this paper, we relax the requirement of gold-standard human-validated PROLEX and transcripts, by leveraging the LanguageNet G2P models [25]. The LanguageNet G2P models are available for 142 languages. The phonemization procedure in Whistle is detailed in Section IV-B.
For crosslingual speech recognition, denote the phoneme inventory for a new, target language (unseen in pre-training) by $V_{\text{cross}}$. For recognizing speech from the target language, we can initialize a CTC-based model from the pre-trained encoder. The embeddings corresponding to the phonemes in $V_{\text{multi}}\cap V_{\text{cross}}$ are directly copied for initialization. For those phonemes that are not included in the multilingual phoneme alphabet $V_{\text{multi}}$ but appeared in the target language inventory $V_{\text{cross}}$, we randomly initialize their phoneme embeddings. The initialized CTC model can then be fine-tuned over labeled speech from the target language. In this way, the fine-tuned encoder and phoneme embeddings can be used to calculate the logits and the posteriori distribution of $\pi_{t}$ in CTC, and WFST-based decoding can be applied for recognizing speech from the target language.
III-B Subword-based multilingual supervised pre-training
Multilingual supervised pre-training based on subwords is very similar to that based on phonemes, as described in Section III-A, which can still base on the CTC method and use WFST-based decoding with word-level n-gram language model. The major difference is that subword-based multilingual supervised pre-training employs subwords for labeling. Thus, the alphabet of labels $V$ consists of subwords; the lexicon for WFST-based decoing is an orthography lexicon (i.e., words are formed by a sequence of subwords); The row vectors from the matrix $W$ could be viewed as embeddings for subwords. Converting text into subwords is often referred to tokenization, which is still under investigation and needs a balance between granularity and ASR performance [12].
In this paper, we use Byte Pair Encoding (BPE) based subwords, or say, tokens [48]. BPE introduces a word segmentation algorithm, which initializes the token alphabet with the character alphabet and iteratively merges the most frequent pair of tokens. In this way, BPE obtains a compact token vocabulary of variable-length subword units. Notably, the merging of tokens in BPE is based on their frequencies. A straightforward application of BPE may inappropriately favor the merging from high-resource languages; for low-resource languages, tokens may be mostly single characters. Similar to [49], sentences are sampled according to a multinomial distribution with probabilities $\left\{q_{l}\right\}_{l=1...N}$:
$q_{l}=\frac{p_{l}^{\beta}}{\sum_{i=1}^{N}p_{i}^{\beta}}\quad with\quad p_{l}=\frac{n_{l}}{\sum_{i=1}^{N}n_{i}},$ (2)
where $\beta$ controls the sampling of languages with different frequencies. We use $\beta$ = 0.5 in experiments. $N$ is the number of seen languages in the training data, and $n_{l}$ denotes the number of sentences for language $l$. By such data sampling, we can increase the number of tokens associated to low-resource languages and reduce the bias towards high-resource languages.
III-C Multilingual self-supervised pre-training
We pre-train a wav2vec 2.0 model [28] on our multilingual pre-training data (just audio data). The basic architecture of the wav2vec 2.0 model is as follows. A convolutional feature encoder maps raw audio $x_{1:T}$ to latent speech features $z_{1},\dots,z_{T}$, which are then fed to a Transformer to output contextual representations $h_{1},\dots,h_{T}$ [50, 37]. The Transformer architecture is the same as in BERT [51, 37]. During training, a quantization module is employed to discretize the latent features $z_{1},\dots,z_{T}$ to $q_{1},\dots,q_{T}$, which represent the targets in the contrastive learning objective. The quantization module uses a Gumbel softmax to choose entries from the codebooks and the chosen entries are concatenated to be $q_{1},\dots,q_{T}$ [52, 53, 50]. The wav2vec 2.0 model is trained by solving a contrastive task on masked feature encoder outputs. During training, spans of ten time steps with random starting indices are masked. The objective is to predict the true quantized latent $q_{t}$ for masked time-steps within a set of $K=100$ distractors sampled from other masked time steps.
Basically, the pre-trained wav2vec 2.0 model is only an acoustic encoder, consisting of a convolutional feature encoder and a transformer contextual encoder. In order to recognize speech from any language, we need to introduce a linear layer (parameterized by matrix $W$) followed by softmax on top of the encoder output $h_{1},\cdots,h_{T}$, as shown in Eq. (1), and perform fine-tuning over labeled data. The labels could be in the form of either phonemes or subwords.
IV Experimental setup
IV-A Dataset
TABLE I: Multilingual and crosslingual data information, including the language code, the language family, the number of IPA phonemes, the sizes of train, development and test sets in hours.
Code Language Family IPA Hours
Train Dev Test
Multi. en English West Germanic 39 2227.3 27.2 27.0
es Spanish Romance 32 382.3 26.0 26.5
fr French Romance 33 823.4 25.0 25.4
it Italian Romance 30 271.5 24.7 26.0
ky Kyrgyz Turkic 32 32.7 2.1 2.2
nl Dutch West Germanic 39 70.2 13.8 13.9
ru Russian East Slavic 32 149.8 14.6 15.0
sv Swedish North Germanic 33 29.8 5.5 6.2
tr Turkish Turkic 41 61.5 10.1 11.4
tt Tatar Turkic 31 20.8 3.0 5.7
Cross. pl Polish West Slavic 35 129.9 11.4 11.5
id Indonesian Austronesian 35 20.8 3.7 4.1
We conduct experiments on the CommonVoice dataset [54] released at September 2022 (v11.0). CommonVoice is a large multilingual speech corpus, with spoken content taken primarily from Wikipedia articles. We select ten languages for multilingual pre-training experiments: English (en), Spanish (es), French (fr), Italian (it), Kyrgyz (ky), Dutch (nl), Russian (ru), Swedish (sv), Turkish (tr) and Tatar (tt), with a total of 4069.3 hours, which cover rich language families. We refer to this dataset of 10 languages as CV-Lang10. We select Polish (pl) and Indonesian (id) for crosslingual finetuning experiments, which are from two unseen language families. Detailed database descriptions are shown in Table I. We combine all data from the ten languages to form the training, development, and test sets for multilingual pre-training experiments. For each language, we use its transcripts to separately train a word-level n-gram language model for WFST-based decoding.
TABLE II: Example transcriptions for each language in CV-Lang10
Code Text transcript Transcription with subwords Transcription with IPA symbols
[…]
es no lo he visto no lo he v ist o n o l o e b i s t o
fr vous ne me comprenez pas vous ne me comp ren ez pas v y n m k  p    n e p a
it è meglio separarci adesso è me g lio separ ar ci ad esso m e  i o s e p a r a r t   a r s s o
ky менин эч кандай кuнм жок мен ин эч кандай кuн м жок m e n i n e t k  n d  j k y n ø m d   o k
nl ze is een bekend model ze is een bek end mod el z e  s e n b  k  n t m o d  l
ru база данных обновлена ба за дан ных об нов лен а b a z a d a n  x o b n o v l e n a
[…]
tr bunlar en büyükleri bun lar en b üy ük leri b u n  a  e n b y j y k l e r i
tt меншулай яшп ятабыз мен шулай яш п я та быз m j e n æ  u l a j j a  æ p j a t a b  z
pl lubię muzykę klasyczną lu b ię mu zy kę k la sy cz ną l u a b v i  m u w z  k  k l a t s   n
id semoga cepat sembuh sem o ga c ep at sem b uh s  m   a t   p a t s  m b o h
IV-B Text normalization and phonemization
For text normalization, all punctuation marks are removed, except those that affect pronunciation (such as the apostrophe in English). Certain sentences contain many foreign words are discarded, since G2P converters cannot properly convert them. For reproducible research, details of text normalization and the IDs of deleted sentences for each language will be released in our public repository.
The FST (Finite State Transducer) based G2P toolkit, Phonetisaurus [26], is utilized to generate labeling of utterances in IPA phonemes from text transcripts. The trained FSTs for use with Phonetisaurus can be obtained from LanguageNet [25]. Examples of phoneme annotations for each language in CV-Lang10 are shown in Table II. By applying Phonetisaurus G2P tool with LanguageNet FSTs, we can also create a PROLEX for each language, which is needed for WFST-based decoding with phoneme-based CTC model. The phonetic transcripts and the PROLEXs for CV-Lang10 will be released in our public repository.
Remarkably, our phonemization procedure produces weakly phonetic supervision for model training. The FST-based G2P procedure by LanguageNet and Phonetisaurus is not perfect. As noted in [25], PERs ranging from 7% to 45%. We only correct a few obvious labeling errors, but the phoneme labels are still somewhat noisy in general. Additionally, we remove the diacritics and suprasegmentals (like stress and tone) that may be necessary for representing phones, and mainly use base phonemes in our annotation [^2] From phonetics and phonology [19], while phones represent physical speech sounds (and thus language-independent), phonemes are not physical sounds; they are abstract mental representations of the phonological units of a language, the units used to represent words in our mental lexicon (and thus language dependent). A particular realization (pronunciation) of a phoneme is called a phone. The collection of phones that are the realizations of the same phonemes are called the allophones of that phoneme. Phonemes for annotation are thus in a coarser granularity than phones, which may facilitate sharing between languages. The 12 languages examined in this paper are all non-tonal languages. So we preliminarily sidestep the problem how tones should be incorporated in phoneme-based multilingual models. This is a interesting future work, as previously investigated in [55]. . While some recent studies pursue universal phone recognition [14, 31], this paper does not aim for phone recognition. On the one hand, accurate gold-standard phone labeling is hard to obtain. On the other hand, when we use WFST-based decoding with PROLEXs and aim for reducing word error rates (WERs), the complexity of constructing an allophone layer to transform the language-independent phone distributions to the language-dependent distributions may not be necessary. Training with weakly phonetic supervision and decoding with PROLEXs, with phonemes serving as an interface between acoustics and text, is found to obtain superior results in MCL-ASR in our experiments. Presumably, as long as the PROLEXs and the phonetic transcriptions are aligned in some way, weakly phonetic supervision can well drive model learning.
Figure 2: Counts of phonemes (top) and subwords (down) in the CV-Lang10 training set.
IV-C Model training
The CAT toolkit [22] is used for training CTC [4] based ASR models in our experiments. Three sizes of acoustic encoders are used in our experiments, all based on Conformer [45] networks. The small-sized Conformer encoder (S) consists of 14 encoder blocks with dimension 512. We set the self-attention layer to have 4 heads with 36-dimension hidden states, and the feed-forward network (FFN) dimension to 512. The middle-sized Conformer encoder (M) uses 22 blocks, model dimension 640, FFN dimension 640, attention dimension 160, while the large-sized Conformer encoder (L) uses 22 blocks, model dimension 1024, FFN dimension 1024, attention dimension 224. For phoneme-based models, the multilingual alphabet size of phonemes is 73. For subword-based models, the multilingual alphabet size of subwords is 4998. Counting statistics for phonemes and subwords over CV-Lang10 are shown in Figure 2.
[…]
TABLE III: Phoneme error rates (PERs) for phoneme-based monolingual models and multilingual pre-trained models on the CV-Lang10 dataset. (S: small, M: middle, L: large)
id Model Size (M) en es fr it ky nl ru sv tr tt Avg.
O1 Mono. phoneme 90 7.39 2.47 4.93 2.87 2.23 4.60 2.72 18.69 6.00 10.54 6.11
M1 Multi. phoneme S 90 8.02 3.37 5.68 4.04 8.29 5.77 6.05 18.07 8.32 8.53 7.61
M2 Multi. phoneme M 218 6.70 2.63 4.53 3.12 5.95 3.95 4.61 14.81 6.04 8.47 6.08
M3 Multi. phoneme L 543 5.42 1.96 3.52 2.25 4.06 2.64 2.97 11.33 4.04 5.97 4.41
TABLE IV: Word error rates (WERs) for phoneme-based monolingual models and multilingual pre-trained models on the CV-Lang10 dataset, compared with the subword-based multilingual pre-trained model.
id Model Size (M) en es fr it ky nl ru sv tr tt Avg.
O1 Mono. phoneme 90 10.59 7.91 15.58 9.26 1.03 8.84 1.62 8.37 8.46 9.75 8.14
M4 Multi. subword 92 12.00 9.82 12.40 9.98 3.29 9.67 3.31 9.95 9.11 13.56 9.30
M1 Multi. phoneme S 90 10.76 8.68 16.01 9.98 1.02 7.32 1.59 6.14 7.63 7.30 7.64
M2 Multi. phoneme M 218 9.83 7.82 14.94 9.04 0.91 6.57 1.65 5.65 7.27 7.37 7.10
M3 Multi. phoneme L 543 8.80 7.02 14.02 8.16 0.94 6.22 1.46 5.06 7.05 6.92 6.56
V Results
In the following, we introduce the experimental results over CV-Lang10, which serves as a common setup for comparing the three MCL-ASR approaches - supervised pre-training with weakly phonetic supervision (Whistle), subword-based supervised pre-training, and wav2vec 2.0 based self-supervised pre-training. The three approaches are described in Section III-A, III-B, and III-C, respectively. An MCL-ASR approach is usually evaluated under two tasks. The first is to recognize utterances from seen languages, i.e., the languages that are included in multilingual pre-training. The second is to recognize utterances from unseen languages, i.e., crosslingual speech recognition, which is often performed by fine-tuning the model obtained from pre-training.
V-A Multilingual pre-training
On the CV-Lang10 dataset, 10 phoneme-based monolingual models are trained, each for a single language and with 90M parameters. Phoneme-based multilingual models (Whistle models) and subword-based multilingual models are trained for comparison. WFST-based decoding are used for all models. The PERs and WERs are shown in Table III and IV respectively. The main observations are as follows.
First, comparing within phoneme-based models, it can be seen that pooling data from multiple languages and training multilingual models clearly reduces PERs over monolingual models, as shown in prior works [14, 31]. Particularly, a single multilingual model (Mult. phoneme L with 543M parameters) performs significantly better than the 10 monolingual separately-trained models (10 * 90M parameters), on averaged PERs over the 10 seen languages. Furthermore, we can see that reductions in WERs can be obtained as well, by phoneme-based multilingual pre-training and WFST-based decoding. Interestingly, in terms of WERs, even the small multilingual model (Mult. phoneme S with 90M parameters) surpasses the monolingual models.
Second, comparing multilingual models based on phonemes and subwords, it is found that the phoneme-based multilingual model (M1) obtains better WERs than the subword-based multilingual model (M4), with close model sizes (around 90M) [^5] The minor difference in model sizes between phoneme-based model and subword-based model (90M vs 92M) is due to the size of the linear layer because of the different alphabet sizes., with 18% WER relative reduction [^6] An exception is that for French, the phoneme-based multilingual model does not outperform the subword-based multilingual model in WER, though the PERs are good. From the statistics of CV-Lang10, we find that the percentage of homophones in the G2P PROLEX of French is the highest (22.5%). The other large percentages of homophones in the 10 langauges in CV-Lang10 is 9.0% for English, 5.2% for Spanish, while others are below 3%. Moreover, it is found that some consonants in French words are usually not pronounced, but they may be pronounced when they are spoken in sentences. The WFST-based decoding with a PROLEX may not be good at capturing these regularities. These issues could be alleviated by developing a better method of decoding from phonemes, which will be explored in future.. This is a fair comparison to answer RQ-1, since both models are trained with the same dataset and the same encoder architecture. Presumably, compared to using subwords which mainly serve for text writing, using phonemes as labels is more natural and better for sound classification, since inherently they are more directly related to describing sounds for languages. Moreover, it can been seen from Figure 2 that data imbalance is more severe in subword supervision than phoneme supervision. From a machine learning perspective, multi-task learning could be severely affected by data imbalance. When data are not well balanced in training, an annoying phenomenon, often observed in subword-based systems, is that high resource languages may suffer from interference and low resource languages may be under-trained, which cause performance degradation [36, 12]. Subword-based systems need special tricks to struggle with data imbalance, such as careful tokenization to appropriately creating the set of tokens [12], human-in-the-loop data mixing in training [36]. In contrast, the superior performances from phoneme-based systems are obtained by training on natural data mixing and adopting the classic IPA symbols that have been matured for describing human sounds for a long time.
Third, we can see clear scaling properties of phoneme-based models - PERs and WERs are consistently reduced for both high-resource and low-resource languages, as the model sizes are increased. Again, remarkably, the performance improvements for different sizes of phoneme-based models are obtained by training on natural data mixing.
TABLE V: PERs and WERs for phoneme-based crosslingual fine-tuning (FT) on Polish. The pre-training dataset is CV-Lang10.
id Model 1h 10h 130h
PER WER PER WER PER WER
O2 Mono. phoneme 86.01 99.98 30.38 13.86 2.82 4.97
M5 W2V (En) phoneme FT 25.76 11.09 16.64  6.75 5.80 4.57
M6 W2V (10 lang) phoneme FT 21.10  7.94 12.65  5.65 6.08 4.44
M7 M1 + phoneme FT 17.96  6.95 10.47  5.27 1.97 4.30
TABLE VI: WERs for subword-based crosslingual fine-tuning (FT) on Polish. The pre-training dataset is CV-Lang10.
id Model 1h 10h 130h
w/o  w w/o  w w/o  w
LM LM LM LM LM LM
O3 Mono. subword 98.41 98.38 90.98 59.43 19.38 7.12
M8 W2V (En) subword FT 100 100 45.64 7.08 8.53 3.85
M9 W2V (10 lang) subword FT 99.97 100 36.93 5.71 7.49 3.45
M10 M4 + subword FT 70.13 9.16 31.90 4.89 5.44 3.76
M11 M1 + subword FT 69.50 8.63 31.89 4.83 5.84 3.82
TABLE VII: PERs and WERs for phoneme-based crosslingual fine-tuning (FT) on Indonesian. The pre-training dataset is CV-Lang10.
id Model 1h 10h 20h
PER WER PER WER PER WER
O4 Mono. phoneme 96.52 100 27.30 7.71 5.74 3.28
M12 W2V (En) phoneme FT 31.30 6.73 10.89 3.31 6.84 2.83
M13 W2V (10 lang) phoneme FT 24.91 3.75 10.32 2.79 6.30 2.47
M14 M1 + phoneme FT 21.64 3.27 7.90 2.54 4.79 2.43
TABLE VIII: WERs for subword-based crosslingual fine-tuning (FT) on Indonesian. The pre-training dataset is CV-Lang10.
id Model 1h 10h 20h
w/o  w w/o  w w/o  w
LM LM LM LM LM LM
O5 Mono. subword 96.62 96.42 69.57 49.67 31.96 10.85
M15 W2V (En) subword FT 100 100 19.98 5.28 11.68 3.59
M16 W2V (10 lang) subword FT 99.64 99.97 19.08 4.52 12.01 3.15
M17 M4 + subword FT 64.00 23.56 19.41 3.91 13.15 3.07
M18 M1 + subword FT 67.71 24.57 18.21 3.59 12.48 2.92
V-B Crosslingual fine-tuning
Over the CV-Lang10 dataset, we obtain the phoneme-based supervised pre-trained model (M1), which can be further fine-tuned with either phoneme labels or subword labels for crosslingual speech recognition. The subword-based supervised pre-trained model (M4) is fine-tuned with subword labels for crosslingual speech recognition. The wav2vec 2.0 models, “W2V (10 lang)” and “W2V (En)”, can be fine-tuned with either phoneme labels or subword labels for crosslingual speech recognition. The four pre-trained models used in the crosslingual experiments all have the same model size (around 90M parameters). On the four pre-trained models, we perform full-parameter fine-tuning, except that for the two wav2vec 2.0 based pre-trained models, the convolutional feature encoder are frozen.
To test different multilingual pre-trained models for crosslingual speech recognition, we conduct phoneme-based and subword-based crosslingual fine-tuning on unseen languages. The training data from an unseen language is divided into three scales to simulate different resource scenarios, while the test and validation data remain unchanged.
The first unseen language is Polish. Polish has 31 phonemes contained in CV-Lang10 and 4 unseen phonemes. The training data is divided into three scales: 1 hour, 10 hours, and full (130 hours). Combining Table V and Table VI, we have the following main observations.
  • In the low-resource scenario with 1-hour Polish training data, phoneme pre-training (PT) followed by phoneme fine-tuning (FT) performs the best (6.95). Results with phoneme PT are much better than those with subword PT, which clearly shows the advantage of phonetic supervision in representation learning from multilingual data (RQ-1). When comparing phoneme PT and wav2vec 2.0 PT (M7 vs M6, M11 vs M9), phoneme PT shows obvious superiority (RQ-2).
  • In the scenario with 10-hour Polish training data, the performance with subword PT models begins to improve. When followed by subword FT, both phoneme PT and subword-based PT show equally excellent results (4.83 and 4.89).
  • With the full Polish training data, the wav2vec 2.0 PT models start to perform well, surpassing results with both subword PT and phoneme PT (3.45 ¡ 3.76 ¡ 3.82). This may reflect some benefit of wav2vec 2.0 PT when fine-tuned with abundant labels, but such top-performing result with wav2vec 2.0 PT is not observed in Indonesian experiments, as shown below.
The second unseen language is Indonesian. All 35 phonemes of Indonesian are contained in CV-Lang10. But Indonesian belongs to the Austronesian language family, which are somewhat more different from CV-Lang10, and only 20 hours of training data are available. These make crosslingual fine-tuning for Indonesian more challenging. The training data is divided into three scales: 1 hour, 10 hours, and full (20 hours).
From Table VII and Table VIII for Indonesian, the observations are similar to those for Polish. In the more challenging scenario with larger linguistic difference and less training data, the advantages of phoneme PT followed by phoneme FT are more obvious, across all the three scales of data settings. It seems that when training data are more limited, the better results can be obtained by phoneme supervision, compared to subword supervision and self-supervision. When the amount of crosslingual training data increases, the performance gaps between phoneme supervision, subword supervision and self-supervision may diminish. Presumably, the fine-tuning with abundant data behaves like end-to-end monolingual training and the effect of different PT methods may become weak.
Figure 3: Visualization of embeddings by t-SNE. (a) Phoneme embeddings from M1, (b) Subword embeddings from M4. In (a), blue indicate the consonants and red indicate the vowels.
VI Ablation study
VI-A Analysis of embeddings
To gain intuitive understanding of the multilingual models trained under phonetic supervision and graphemic supervision, we apply t-SNE [57] to draw the 512-dimensional embeddings on a 2-dimensional map. Figure 3(a) and (b) show the maps of the 73 phoneme embeddings and the 4998 subword embeddings, obtained from the phoneme-based model M1 and subword-based model M4, respectively. By comparing the two figures, it can be easily seen that the phoneme embeddings are more evenly dispersed in the high-dimensional space. In contrast, subword embedings are densely crowded in the center and become sparser as they move outward. This indicates that the representation learning in the subword-based model is not so balanced as in the phoneme-based model. Presumably, this is due to the severe data imbalance in subword supervision. Furthermore, it can be noticed that most of the vowels embeddings cluster in the bottom right area of Figure 3(a). Certain consonant phonemes, like approximants (”, ” and ’j’), also appear in this region, since approximants fall between fricatives and vowels. This reflects that the phoneme-based model not only learns the differences between phonemes, but also captures some phonetic similarities between phonemes.
VI-B Test of catastrophic forgetting
In previous sections, we show the advantage of multilingual pre-trained models by phoneme supervision over those by subword supervision for recognizing seen and unseen languages. We see that after a pre-trained multilingual model is fine-tuned over data from a new language, the fine-tuned multilingual model can recognize speech from the new language. Then, to what degree the performance of the fine-tuned multilingual model on previous seen languages would be affected? This is an interesting question for continual pre-training of multilingual models to support more new languages, a question related to catastrophic forgetting of neural network based models [58]. A complete investigation into continual pre-training of multilingual models is outside the scope of this paper. Here we present a preliminary examination of the two approaches, phoneme or subword-based multilingual models, in overcoming catastrophic forgetting.
[…]
TABLE IX: Test of catastrophic forgetting for the multilingual models, pre-trained over CV-Lang10 and fine-tuned on 10 minutes of a new language (Polish). WARD denotes word accuracy relative degradation of the averaged WER over the ten old languages in CV-Lang10.
id Model pl en fr es it ru nl tr ky sv tt Avg. WARD
M19 M1 + 10min phoneme FT 11.0 68.5 69.3 57.1 50.3 48.3 60.9 31.8 58.4 42.3 33.0 52.0 48
M20 M4 + 10min subword FT 93.2 92.2 95.0 92.5 92.5 262.5 103.6 241.5 125.9 180.5 254.4 154.1 160
VI-C Training efficiency
[…]
TABLE X: Training efficiency of phoneme-based and subword-based pre-training (PT) and fine-tuning (FT).
id Model Batch size Epochs for converging
M1 phoneme PT 640 63
M11 M1 + pl subword FT 320 195
M4 subword PT 640 83
M10 M4 + pl subword FT 320 223
VII Conclusions and future work
This paper starts from examining the pros and cons of the three main approaches for MCL-ASR - supervised pre-training with phonetic transcription or graphemic transcription, and self-supervised pre-training. We find that pre-training with phonetic supervision has been underappreciated so far for MCL-ASR, while conceptually it is more advantageous for information sharing between different languages. This paper explores the approach of pre-training with weakly phonetic supervision towards data-efficient MCL-ASR, which is called Whistle. We relax the requirement of gold-standard human-validated phonetic transcripts, and obtain IPA based transcripts by leveraging Phonetisaurus (an FST based G2P toolkit) with LanguageNet G2P FSTs. We construct a common experimental setup based on the CommonVoice dataset, called CV-Lang10, with 10 seen languages and 2 unseen languages (Polish and Indonesian). A set of experiments are conducted on CV-Lang10 to compare, as fair as possible, the three approaches under the common setup for MCL-ASR. Training with weakly phonetic supervision (though somewhat noisy) and decoding with PROLEXs, with phonemes serving as an interface between acoustics and text, is found to obtain superior results in MCL-ASR in our experiments, in terms of speech recognition for seen languages, crosslingual performance for unseen languages with different amounts of few-shot data, overcoming catastrophic forgetting, and training efficiency. Moreover, phoneme-based models naturally overcome language imbalance and can be efficiently trained on natural data mixing, while subword-based models need careful tokenization and data mixing in training. When training data is more limited, phoneme supervision can achieve better results compared to subword supervision and self-supervision, thereby providing higher data-efficiency.
This work demonstrates some advantages of weakly phonetic supervision towards data-efficient MCL-ASR. There are interesting directions for future work. First, we preliminarily sidestep the problem how tones should be incorporated in pre-training multilingual phoneme-based models, since the 12 languages examined in this paper are all non-tonal languages. There have been some effort towards addressing this problem [55]. Second, this work mainly uses WFST based decoding with PROLEXs. Better methods of decoding from phonemes could be explored in future, such as based on sequence-to-sequence models [59]. Third, scaling the approach of Whistle with more languages and more data is expected to achieve increasingly better MCL-ASR performance. Meanwhile, it is worthwhile to investigate how to incrementally learn from new languages with a non-stationary stream Continual learning methods such as based on prompt pool [60, 61] could be incorporated into MCL-ASR.